OpenStudy (anonymous):

Write an appropriate direct variation equation if y = 8 when x = 4

jimthompson5910 (jim_thompson5910):

y = k*x 8 = k*4 ... plugin x = 4 and y = 8 solve for k

OpenStudy (anonymous):

i dont understand it. but heres the multiple choice A. y = 2x B. y = 32x C. y = -2x D. y = -32x

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

if 8 = k*4 then what is k?

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

so y = k*x turns into y = 2x

jimthompson5910 (jim_thompson5910):

and that's your equation